### v1

# Comment on “Electric conductivity of graphene: Kubo model versus a nonlocal quantum field theory model” (arXiv:2403.02279v3)

Abstract
Recently, Rodriguez-Lopez, Wang, and Antezza [arXiv:2403.02279v3; Phys. Rev. B ${\bf 111}$, 115428 (2025)] compared the theoretical descriptions of electric conductivity of graphene given by the Kubo model and quantum field theory in terms of the polarization tensor. According to these authors, in the spatially nonlocal case, the quantum field theoretical description contains “hard inconsistencies”. To bring the predictions of quantum field theory in agreement with those following from the Kubo model, the modified expression was used which relates the conductivity and the polarization tensor. Here, it is shown that this modification violates the requirement of gauge invariance and, thus, is unacceptable. By comparing both theoretical approaches, we demonstrate that all the results obtained within quantum field theory are physically well justified whereas an application of the modified expression for the conductivity of graphene leads to the consequences of nonphysical character.
Reference 1 compares the expressions for the electric conductivity of graphene. One of them was obtained 2 ; 3 ; 4 ; 5 in the framework of Dirac model using the formalism of quantum field theory and, specifically, the concept of the polarization tensor 6 ; 7 ; 8 ; 9 ; 10 . The other one was found using the nonlocal Kubo model 1 . According to Ref. 1 , in the spatially nonlocal region, the transverse conductivity of graphene obtained using quantum field theory shows a “nonphysical plasmalike behavior… at low frequencies” leading to “dissipation-less permanent currents”. To bring the results obtained using quantum field theory in agreement with those found using the nonlocal Kubo model, Ref. 1 modifies the expression connecting the polarization tensor with the tensor of electric conductivity.
Below we demonstrate that the modified expression used in Ref. 1 is in violation of the requirement of gauge invariance. It is shown that the results found using this expression are physically unacceptable. Thus, no modification is needed, and the claims of Ref. 1 against the quantum field theoretical description are invalid. We also list several other inconsistencies contained in Ref. 1 .
[…]
In the absence of constant magnetic field, the polarization tensor of graphene can be expressed via two independent quantities, e.g. 1 ; 2 ; 3 ; 4 ; 5 ; 13 ; 14 , ${}\Pi_{00}(\omega,\mbox{\boldmath$q$})$ and
$\Pi(\omega,\mbox{\boldmath$q$})\equiv q^{2}\Pi_{\mu}^{\,\mu}(\omega,\mbox{\boldmath$q$})+\left(\frac{\omega^{2}}{c^{2}}-q^{2}\right)\,\Pi_{00}(\omega,\mbox{\boldmath$q$}).$ (1)
In terms of these quantities, it takes the form 7
$\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})=\left(\begin{array}[]{ccc}\Pi_{00}&\frac{q_{0}q_{1}}{q^{2}}\Pi_{00}&\frac{q_{0}q_{2}}{q^{2}}\Pi_{00}\\[8.53581pt] \frac{q_{0}q_{1}}{q^{2}}\Pi_{00}&\frac{q_{0}^{2}q_{1}^{2}}{q^{4}}\Pi_{00}-\frac{q_{2}^{2}}{q^{4}}\Pi&\frac{q_{1}q_{2}}{q^{4}}(q_{0}^{2}\Pi_{00}+\Pi)\\[8.53581pt] \frac{q_{0}q_{2}}{q^{2}}\Pi_{00}&\frac{q_{1}q_{2}}{q^{4}}(q_{0}^{2}\Pi_{00}+\Pi)&\frac{q_{0}^{2}q_{2}^{2}}{q^{4}}\Pi_{00}-\frac{q_{1}^{2}}{q^{4}}\Pi\end{array}\right)$ (2)
where $q^{2}=q_{1}^{2}+q_{2}^{2}$ and $q_{0}=\omega/c$.
According to Ref. 1 , the tensor of electric conductivity used in quantum field theoretical formalism of Refs. 2 ; 3 ; 4 ; 5 ; 10 ; 12 ; 18 ; 19 is given by $\sigma_{\mu\nu}=\Pi_{\mu\nu}/(-i\omega)$. In fact, however, Refs. 10 ; 12 ; 19 use this equation with $+i\omega$ in denominator. As to Refs. 2 ; 3 ; 4 ; 5 ; 18 , they consider only the longitudinal and transverse conductivities. It is pertinent to note that Ref. 1 does not specify the used physical units and throughout the paper presents equations written in different systems of units (see below). Here, we present all the following equations in the Gaussian units. For the sake of clarity, we preserve the velocity of light $c$, the Planck constant $\hbar$, and the Boltzmann constant $k_{B}$.
Under this convention, the electric conductivity tensor used in Refs. 10 ; 12 ; 15 ; 16 ; 19 is given by
$\sigma_{\mu\nu}(\omega,\mbox{\boldmath$q$})=\frac{c^{2}}{4\pi\hbar}\,\frac{\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})}{i\omega}.$ (3)
As discussed above, Ref. 1 considers Eq. (3) as unsatisfactory and writes it in the modified “regularized” form:
$\sigma_{\mu\nu}^{\rm K}(\omega,\mbox{\boldmath$q$})$ $=$ $\frac{c^{2}}{4\pi\hbar}\,\frac{\widetilde{\Pi}_{\mu\nu}(\omega,\mbox{\boldmath$q$})}{i\omega}$ (4)
$\equiv$ $\frac{c^{2}}{4\pi\hbar}\,\frac{\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})-\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})}{i\omega}.$
Note that Eq. (4) differs from the corresponding Eqs. (3), (93), and (95) in Ref. 1 by the constant factor which is not important in the given context.
From Eq. (2) one obtains
$\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})=\left(\begin{array}[]{ccc}\lim\limits_{\omega\to 0}\Pi_{00}&0&0\\ 0&-\frac{q_{2}^{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi&\frac{q_{1}q_{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi\\ 0&\frac{q_{1}q_{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi&-\frac{q_{1}^{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi\end{array}\right).$ (5)
Due to the gauge invariance, the polarization tensor $\Pi_{\mu\nu}$ satisfies the transversality condition $q_{\mu}\Pi^{\mu\nu}(\omega,\mbox{\boldmath$q$})=0$. Due to Eq. (3), the same condition $q_{\mu}\sigma^{\mu\nu}(\omega,\mbox{\boldmath$q$})=0$ is also valid for the components of the tensor of electric conductivity.
It is easily seen, however, that the modified polarization tensor $\widetilde{\Pi}_{\mu\nu}$ and the conductivity tensor $\sigma_{\mu\nu}^{\rm K}$ defined in Eq. (4) do not satisfy the transversality condition. Really, from Eq. (5) one finds that
$q_{\mu}\widetilde{\Pi}^{\mu 0}(\omega,\mbox{\boldmath$q$})=-q_{\mu}\lim\limits_{\omega\to 0}\Pi^{\mu 0}(\omega,\mbox{\boldmath$q$})=-q_{0}\lim\limits_{\omega\to 0}\Pi^{00}(\omega,\mbox{\boldmath$q$})\neq 0$ (6)
because even for the sheet of pristine graphene ($\Delta=\mu=0$) at fixed $q$ it holds 21
$\lim\limits_{\omega\to 0}\Pi^{00}(\omega,\mbox{\boldmath$q$})=\frac{\alpha\pi\hbar cq}{v_{F}}+\frac{16\alpha c}{v_{F}^{2}}k_{B}T\ln 2,$ (7)
where $\alpha=e^{2}/(\hbar c)$ is the fine structure constant and $v_{F}\approx c/300$ is the Fermi velocity of graphene.
Using Eqs. (4), (6) and (7), we find
$q_{\mu}\sigma^{{\rm K},\mu 0}(\omega,\mbox{\boldmath$q$})=-i\frac{\alpha c^{2}}{4\pi\hbar v_{F}}\left(\pi\hbar q+\frac{16k_{B}T}{v_{F}}\ln 2\right)\neq 0,$ (8)
i.e., for the modified polarization and conductivity tensors used in Ref. 1 the transversality condition for $\nu=0$ is violated.
[…]
$\Pi_{ij}(\omega,\mbox{\boldmath$q$})=\frac{q_{i}q_{j}}{q^{2}}\Pi_{L}(\omega,\mbox{\boldmath$q$})+\left(\delta_{ij}-\frac{q_{i}q_{j}}{q^{2}}\right)\Pi_{T}(\omega,\mbox{\boldmath$q$}),$
$\sigma_{ij}(\omega,\mbox{\boldmath$q$})=\frac{q_{i}q_{j}}{q^{2}}\sigma_{L}(\omega,\mbox{\boldmath$q$})+\left(\delta_{ij}-\frac{q_{i}q_{j}}{q^{2}}\right)\sigma_{T}(\omega,\mbox{\boldmath$q$}).$ (9)
We set the components of the first equality in Eq. (9) equal to the corresponding components in Eq. (2) and obtain
$\Pi_{L}(\omega,\mbox{\boldmath$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\Pi_{00}(\omega,\mbox{\boldmath$q$}),\quad\Pi_{T}(\omega,\mbox{\boldmath$q$})=-\frac{1}{q^{2}}\Pi(\omega,\mbox{\boldmath$q$}).$ (10)
In a similar way, using the conductivity tensor (3) and the second equality in Eq. (9), one finds
$\sigma_{L}(\omega,\mbox{\boldmath$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\sigma_{00}(\omega,\mbox{\boldmath$q$}),\quad\sigma_{T}(\omega,\mbox{\boldmath$q$})=-\frac{1}{q^{2}}\sigma(\omega,\mbox{\boldmath$q$}),$ (11)
where
$\sigma(\omega,\mbox{\boldmath$q$})=q^{2}\sigma_{\mu}^{\,\mu}(\omega,\mbox{\boldmath$q$})+\left(\frac{\omega^{2}}{c^{2}}-q^{2}\right)\sigma_{00}(\omega,\mbox{\boldmath$q$})=\frac{c^{2}}{4\pi\hbar i\omega}\Pi(\omega,\mbox{\boldmath$q$}).$ (12)
By representing the modified polarization and conductivity tensors (4) used in Ref. 1 in the form of Eq. (9) and repeating the same calculations with the help of Eqs. (2) and (5), we arrive at
$\widetilde{\Pi}_{L}(\omega,\mbox{\boldmath$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\Pi_{00}(\omega,\mbox{\boldmath$q$})=\Pi_{L}(\omega,\mbox{\boldmath$q$}),$
$\widetilde{\Pi}_{T}(\omega,\mbox{\boldmath$q$})=-\frac{1}{q^{2}}\left[\Pi(\omega,\mbox{\boldmath$q$})-\lim\limits_{\omega\to 0}\Pi(\omega,\mbox{\boldmath$q$})\right]$ (13)
and
${\sigma}_{L}^{\rm K}(\omega,\mbox{\boldmath$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\sigma_{00}(\omega,\mbox{\boldmath$q$})=\sigma_{L}(\omega,\mbox{\boldmath$q$}),$
$\sigma_{T}^{\rm K}(\omega,\mbox{\boldmath$q$})=-\frac{1}{q^{2}}\left[\sigma(\omega,\mbox{\boldmath$q$})-\lim\limits_{\omega\to 0}\sigma(\omega,\mbox{\boldmath$q$})\right]$
$=\frac{ic^{2}}{4\pi\hbar\omega q^{2}}\left[\Pi(\omega,\mbox{\boldmath$q$})-\lim\limits_{\omega\to 0}\Pi(\omega,\mbox{\boldmath$q$})\right].$ (14)
As correctly concluded in Ref. 1 , the made modification (4) does not change the longitudinal conductivity of graphene and modifies only the transverse one. Below we focus on the transverse part and demonstrate that the results obtained using this modification are physically unacceptable.
Thus, for a pristine graphene at $T=0$ the quantity $\Pi$ is given by 2 ; 3 ; 4 ; 5 ; 13 ; 14
$\Pi(\omega,\mbox{\boldmath$q$})=\frac{\pi e^{2}q^{2}}{c^{2}}\left\{\begin{array}[]{ll}\phantom{\mp i}\sqrt{v_{F}^{2}q^{2}-\omega^{2}},&|\omega|<v_{F}q,\\[8.53581pt] \mp i\sqrt{\omega^{2}-v_{F}^{2}q^{2}},&|\omega|>v_{F}q,\end{array}\right.$ (15)
where here and below the upper and lower signs are for $\omega>v_{F}q$ and $\omega<-v_{F}q$, respectively.
From Eq. (15) at fixed $q$ one finds
$\lim\limits_{\omega\to 0}\Pi(\omega,\mbox{\boldmath$q$})=\frac{\pi e^{2}v_{F}q^{3}}{c^{2}}.$ (16)
This result is valid not only at $T=0$ but at any temperature 21 . Using Eqs. (14)-(16), for the modified transverse conductivity we obtain
$\sigma_{T}^{\rm K}(\omega,\mbox{\boldmath$q$})=\frac{\sigma_{0}}{\omega}\left\{\begin{array}[]{ll}i\left(\!\sqrt{v_{F}^{2}q^{2}-\omega^{2}}-v_{F}q\right),&|\omega|<v_{F}q,\\[8.53581pt] \pm\sqrt{\omega^{2}-v_{F}^{2}q^{2}}-iv_{F}q,&|\omega|>v_{F}q,\end{array}\right.$ (17)
where $\sigma_{0}=e^{2}/(4\hbar)$ is the universal conductivity of graphene. Note that the regions $0<\omega<v_{F}q$ and $v_{F}q<\omega<cq$ correspond to the evanescent waves whereas for the propagating waves $0\leqslant cq\leqslant\omega$ holds.
As is seen in Eq. (17), for $\omega\geqslant cq$ Ref. 1 arrives at
${\rm Im}\,\sigma_{T}^{\rm K}(\omega,\mbox{\boldmath$q$})=-\frac{\sigma_{0}v_{F}q}{\omega}$ (18)
in contradiction with the previously obtained conclusion that at zero temperature the conductivity of pure graphene in the region of propagating waves is real at all frequencies 23 ; 24 . The result following from the quantum field theoretical definition (3) of the conductivity tensor
$\sigma_{T}(\omega,\mbox{\boldmath$q$})=\frac{\sigma_{0}}{\omega}\left\{\begin{array}[]{ll}i\sqrt{v_{F}^{2}q^{2}-\omega^{2}},&|\omega|<v_{F}q,\\[8.53581pt] \pm\sqrt{\omega^{2}-v_{F}^{2}q^{2}},&|\omega|>v_{F}q\end{array}\right.$ (19)
is in agreement with this conclusion.
According to Ref. 1 , the relation between the polarization tensor and the electric current
$J_{\mu}(\omega,\mbox{\boldmath$q$})=\frac{c}{4\pi h}\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})A^{\nu}(\omega,\mbox{\boldmath$q$})$ (20)
used in Refs. 10 ; 12 ; 15 ; 16 ; 19 for a derivation of Eq. (3) is unsatisfactory because it leads to a nonzero electric current expressed via $\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{\boldmath$q$})$ for the equal to zero electric field. Based on this, Ref. 1 replaces $\Pi_{\mu\nu}$ in Eq. (20) with the modified polarization tensor $\widetilde{\Pi}_{\mu\nu}$ defined in Eq. (4). Below we show that this replacement has no justification.
Really, Ref. 1 uses the temporal gauge and considers the zero electric field, $E^{\nu}(t,\mbox{\boldmath$q$})=0$, as the negative derivative with respect to $t$ of the constant in time vector potential $A_{0}^{\,\nu}(\mbox{\boldmath$q$})$ with fixed $q$. It is true that in the region of evanescent waves, $0<\omega<v_{F}q$, where it is possible to consider limit $\omega\to 0$ at fixed $q$, using Eqs. (2) and (16), one obtains
$\lim\limits_{\omega\to 0}\Pi_{ii}(\omega,\mbox{\boldmath$q$})=-\left(1-\frac{q_{i}^{2}}{q^{2}}\right)\frac{\pi e^{2}v_{F}q}{c^{2}}\neq 0.$ (21)
Thus, in the region of evanescent waves, at any temperature, including $T=0$, the transverse conductivity $\sigma_{T}$ found in the framework of quantum field theory has an imaginary part. For instance, at $T=0$, where it is pure imaginary, using the first line of Eq. (15), we obtain
$\sigma_{T}(\omega,\mbox{\boldmath$q$})=i\sigma_{0}\frac{\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}{\omega}.$ (22)
According to Ref. 1 , Eq. (21) results in the appearance of an electric current in graphene in the absence of electric field, whereas Eq. (22) leads to an unacceptable double pole at zero frequency in the graphene dielectric permittivity. If the expression (4) is used 1 , the modified polarization tensor $\widetilde{\Pi}_{\mu\nu}(\omega,\mbox{\boldmath$q$})$ and the conductivity $\sigma_{T}^{\rm K}(\omega,\mbox{\boldmath$q$})$ vanish in the limit of zero frequency.
Note, however, that at nonzero temperature in the region of propagating waves the pure imaginary current in the absence of electric field arises both in the quantum field theoretical formalism using Eq. (20) and in the approach of Ref. 1 which replaces the polarization tensor $\Pi_{\mu\nu}$ in Eq. (20) with the modified one, $\widetilde{\Pi}_{\mu\nu}$, defined in Eq. (4).
Really, for the propagating waves the condition $0\leqslant cq\leqslant\omega$ holds. Then, if $\omega\to 0$, the wave vector $q$ must go to zero as well. For the real parts of $\Pi_{00}$ and $\Pi$, under the condition $\hbar\omega\ll k_{B}T$, one has 2
${\rm Re}\,\Pi_{00}(\omega,\mbox{\boldmath$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar}\,\frac{q^{2}}{\omega^{2}}\left[1+O\left(\frac{v_{F}^{2}q^{2}}{\omega^{2}}\right)\right],$
${\rm Re}\,\Pi(\omega,\mbox{\boldmath$q$})=8\ln 2\frac{e^{2}k_{B}T}{\hbar}\,\frac{q^{2}}{c^{2}}\left[1+O\left(\frac{v_{F}^{2}q^{2}}{\omega^{2}}\right)\right].$ (23)
With the help of these results, using Eq. (2), we find
${\rm Re}\,\Pi_{11}(\omega,\mbox{\boldmath$q$})={\rm Re}\,\Pi_{22}(\omega,\mbox{\boldmath$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}},$ (24)
i.e., these components of the polarization tensor in the lowest order with respect to the small parameter $v_{F}^{2}q^{2}/\omega^{2}$ are equal to the constant independent on $\omega$ and $q$. Thus, they preserve their value (24) in the limit $\omega,\,q\to 0$.
Using Eqs. (10) and (23), we also obtain
${\rm Re}\,\Pi_{L}(\omega,\mbox{\boldmath$q$})={\rm Re}\,\Pi_{T}(\omega,\mbox{\boldmath$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}.$ (25)
The corresponding values of the longitudinal and transverse conductivities are obtained from Eqs. (3), (11), (12), and (23)
${\rm Im}\,\sigma_{L}(\omega,\mbox{\boldmath$q$})={\rm Im}\,\sigma_{T}(\omega,\mbox{\boldmath$q$})=\sigma_{0}\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega}.$ (26)
For the modified expressions used in Ref. 1 from Eqs. (13), (16), and (23) one finds
${\rm Re}\,\widetilde{\Pi}_{L}(\omega,\mbox{\boldmath$q$})={\rm Re}\,\Pi_{L}(\omega,\mbox{\boldmath$q$}),$
${\rm Re}\,\widetilde{\Pi}_{T}(\omega,\mbox{\boldmath$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}\,+\frac{\pi e^{2}v_{F}q}{c^{2}}.$ (27)
The corresponding conductivities found from Eq. (14) have the imaginary parts
${\rm Im}\,\sigma_{L}^{\rm K}(\omega,\mbox{\boldmath$q$})=\sigma_{0}\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega},$
${\rm Im}\,\sigma_{T}^{\rm K}(\omega,\mbox{\boldmath$q$})=\sigma_{0}\left(\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega}-\frac{v_{F}q}{\omega}\right).$ (28)
What is important, the diagonal components of the modified polarization tensor found from Eq. (9) with added tildes using Eq. (27) do no vanish in the limit $\omega,\,q\to 0$
$\lim\limits_{\omega,\,q\to 0}{\rm Re}\,\widetilde{\Pi}_{11}(\omega,\mbox{\boldmath$q$})=\lim\limits_{\omega,\,q\to 0}{\rm Re}\,\widetilde{\Pi}_{22}(\omega,\mbox{\boldmath$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}.$ (29)
Thus, at nonzero temperature in both the quantum field theory and in the formalism of Ref. 1 a nonzero current in graphene arises even for a zero electric field. Keeping in mind, however, that the conductivity in both cases is pure imaginary because its real part vanishes when $\omega,\,q\to 0$ 2 , this creates no problem. It seems illogical that the same behavior of $\sigma_{L}^{\rm K}$ and $\sigma_{T}^{\rm K}$ in Eq. (28) following from the modified formalism of Ref. 1 in the region of propagating waves as of $\sigma_{L}$ and $\sigma_{T}$ in Eqs. (22) and (26) using quantum field theory is not considered in Ref. 1 as leading to a permanent electric current in the absence of electric field. Note also that just the behavior of the conductivities of graphene according to Eq. (26) leads to the big thermal effect in the Casimir force between two graphene sheets at short separations predicted in Ref. 25 and confirmed experimentally in Refs. 26 ; 27 .
Hence, the modified expression (4) used in Ref. 1 in order to remove the pure imaginary current and the double pole at zero frequency in the transverse dielectric permittivity of graphene in the region of evanescent waves leads to a violation of the condition of gauge invariance and to the physical inconsistencies described above. As to a prediction of the double pole at zero frequency by quantum field theory, it does not contradict to any physical results, including the Kubo approach formulated for the propagating fields, and suggests a path to a resolution of the long-term problems in the Casimir effect 13 .
Last but not least, Ref. 1 does not inform the reader about the used system of units. The inequality in the 9th line of the Abstract contains $\hbar$ and makes a summation of the second power of energy with the second power of mass. From this the reader may conclude that the authors put the speed of light $c$ equal to unity, $c=1$. This conclusion, however, is not supported by the text which explicitly contains the speed of light $c$ in many places [see, e.g., Eq. (18), the third line below Eq. (92), first line below Eq. (114), Eq. (138), first line below Eq. (146), Eqs. (159, (160), first line below Eq. (160), Eq. (168), first line below Eq. (A9), Eqs. (B2), (B7)-(B9), (B11), (B13), (C37), (C40)-(C42), (C51-(C55) etc.]. In spite of this, Ref. 1 makes a summation of the physical quantities of different dimensions in many instances [see, e.g., the last two lines of TABLE I, Eqs. (43), (44), (46), (137), (140), (C1), (C2), line 1 above Eq. (C3) etc.]. In TABLE I, right column, in lines 10 and 11 the components of 3-vectors have different dimensions. In Eqs. (159) and (160) the dimensionless argument of the function $\psi$ is obtained as a ratio of the quantities of different dimensions.
What is more, the expression connecting the polarization and conductivity tensors is written in the form used in Refs. 10 ; 12 where the system of units with $\hbar=c=k_{B}=\varepsilon_{0}=1$ has been used ($\varepsilon_{0}$ is the dielectric permittivity of vacuum). In contradiction with this, Ref. 1 uses the expression for the fine structure constant $\alpha=e^{2}/(\hbar c)$, whereas in Refs. 10 ; 12 the expression $\alpha=e^{2}/(4\pi)$ has been used. The definition of $q_{z}$ below Eq. (120) is in contradiction with the definition of the same quantity given in TABLE I, right column, line 9.
Above Eq. (C34) the authors consider “the case $\Pi_{\mu\nu}(q)=\Pi_{L}(q)$” which is meaningless if to take into account the expression for $\Pi_{L}(q)$ in Eq. (121). The above examples do not exhaust the list of technical inconsistences contained in Ref. 1 .
To conclude, in the foregoing it was shown that the modification of the polarization and conductivity tensors made in Ref. 1 violates their transversality and leads to other physically unacceptable consequences. This modification is superfluous because the claims of Ref. 1 about the problems arising when using the standard formalism of quantum field theory are invalid.
It should be added also that, according to Ref. 1 , the quantum field theoretical description of the conductivity of graphene does not describe unavoidable losses. This is, however, not the case because the conductivity of graphene obtained using this description possesses both the real and imaginary parts. In doing so, the real part of conductivity results in the positive imaginary part of the dielectric permittivity of graphene which describes the losses.
All the above examples were given for the case of a pristine graphene in order to avoid unnecessary difficulties and to make the presentation maximally transparent. However, in the application region of Dirac model, all the statements made remain valid for a general case of graphene possessing some nonzero energy gap and chemical potential.
The authors are grateful to D.V.Vassilevich for useful comments. The work of G.L.K. and V.M.M. was supported by the State Assignment for Basic Research (project FSEG-2023-0016). N.K. was supported in part by the grants 2022/08771-5, 2021/10128-0 of São Paulo Research Foundation (FAPESP).