### v2

# Comment on “Electric conductivity of graphene: Kubo model versus a nonlocal quantum field theory model”

Abstract
Recently, Rodriguez-Lopez, Wang, and Antezza [Phys. Rev. B $\bf 111$}, 115428 (2025)] compared the theoretical descriptions of electric conductivity of graphene given by the Kubo model and quantum field theory in terms of the polarization tensor. According to this article, in the spatially nonlocal case, the quantum field theoretical description contains “hard inconsistencies”. By modifying the equality, which relates the conductivity and polarization expressions, the predictions of quantum field theory were revised and brought in agreement with those following from the nonrelativistic Kubo model. Here, it is shown that this modification violates the requirement of gauge invariance and, thus, is unacceptable. By comparing both theoretical approaches, we demonstrate that all the results obtained within quantum field theory are physically well justified whereas an application of the modified expression for the conductivity of graphene leads to the consequences of nonphysical character.
Reference 1 compares the expressions for the electric conductivity of graphene. One of them was obtained 2 ; 3 ; 4 ; 5 in the framework of Dirac model using the formalism of quantum field theory and, specifically, the concept of the polarization tensor 6 ; 7 ; 8 ; 9 ; 10 . The other one was found using the nonrelativistic nonlocal Kubo model 1 . According to Ref. 1 , in the spatially nonlocal region, the transverse conductivity of graphene obtained using quantum field theory shows a “nonphysical plasmalike behavior… at low frequencies” leading to “dissipation-less permanent currents”. By making a modification of the equality connecting the polarization and electric conductivity expressions, Ref. 1 revised the results obtained using relativistic quantum field theory and brought them in agreement with those found using the nonrelativistic nonlocal Kubo model.
Below we demonstrate that the modified expression used in Ref. 1 is in violation of the requirement of gauge invariance. It is shown that the results found using this expression are physically unacceptable. Thus, no modification is needed, and the objections of Ref. 1 against the quantum field theoretical description are invalid. We also list several other inconsistencies contained in Ref. 1 .
In application to graphene, the formalism of the polarization tensor was elaborated in Refs. 6 ; 7 ; 8 ; 9 and further developed and clarified in Refs. 10 ; 12 ; 13 ; 14 ; 15 ; 16 (see also Ref. 17 and literature therein for the previous publications on this subject). In momentum representation, the components of the polarization tensor $\Pi_{\mu\nu}$ with $\mu,\,\nu=0,\,1,\,2$ depend on the frequency $\omega$ and on the wave vector projection $\mbox{$q$}=(q_{1},\,q_{2})$ on the plane of graphene. For a doped and gapped graphene, they also depend on the energy gap parameter $\Delta$, chemical potential $\mu$ and, at nonzero temperature, on $T$.
In the absence of constant in time, external magnetic field, the polarization tensor of graphene can be expressed via two independent quantities, e.g. 1 ; 2 ; 3 ; 4 ; 5 ; 13 ; 14 , $\Pi_{00}(\omega,\mbox{$q$})$ and
$\Pi(\omega,\mbox{$q$})\equiv q^{2}\Pi_{\mu}^{\,\mu}(\omega,\mbox{$q$})+\left(\frac{\omega^{2}}{c^{2}}-q^{2}\right)\,\Pi_{00}(\omega,\mbox{$q$}).$ (1)
In terms of these quantities, it takes the form 7
$\Pi_{\mu\nu}(\omega,\mbox{$q$})=\left(\begin{array}[]{ccc}\Pi_{00}&\frac{q_{0}q_{1}}{q^{2}}\Pi_{00}&\frac{q_{0}q_{2}}{q^{2}}\Pi_{00}\\[8.53581pt] \frac{q_{0}q_{1}}{q^{2}}\Pi_{00}&\frac{q_{0}^{2}q_{1}^{2}}{q^{4}}\Pi_{00}-\frac{q_{2}^{2}}{q^{4}}\Pi&\frac{q_{1}q_{2}}{q^{4}}(q_{0}^{2}\Pi_{00}+\Pi)\\[8.53581pt] \frac{q_{0}q_{2}}{q^{2}}\Pi_{00}&\frac{q_{1}q_{2}}{q^{4}}(q_{0}^{2}\Pi_{00}+\Pi)&\frac{q_{0}^{2}q_{2}^{2}}{q^{4}}\Pi_{00}-\frac{q_{1}^{2}}{q^{4}}\Pi\end{array}\right),$ (2)
where $q^{2}=q_{1}^{2}+q_{2}^{2}$ and $q_{0}=\omega/c$.
According to Ref. 1 , the tensor of electric conductivity used in quantum field theoretical formalism of Refs. 2 ; 3 ; 4 ; 5 ; 10 ; 12 ; 18 ; 19 is given by $\sigma_{\mu\nu}=\Pi_{\mu\nu}/(-i\omega)$. In fact, however, Refs. 10 ; 12 ; 19 use this equation with $i\omega$ in denominator. As to Refs. 2 ; 3 ; 4 ; 5 ; 18 , they consider only the longitudinal and transverse conductivities. It is pertinent to note that Ref. 1 does not specify the used physical units and throughout the paper presents equations written in different systems of units (see below). Here, we present all the following equations in the Gaussian units. For the sake of clarity, we preserve the velocity of light $c$, the Planck constant $\hbar$, and the Boltzmann constant $k_{B}$.
Under this convention, the electric conductivity tensor used in Refs. 10 ; 12 ; 15 ; 16 ; 19 is given by
$\sigma_{\mu\nu}(\omega,\mbox{$q$})=\frac{c^{2}}{4\pi\hbar}\,\frac{\Pi_{\mu\nu}(\omega,\mbox{$q$})}{i\omega}.$ (3)
As discussed above, Ref. 1 considers Eq. (3) as unsatisfactory and writes it in the modified “regularized” form with the same values of both indices:
$\sigma_{\mu\nu}^{\rm K}(\omega,\mbox{$q$})$ $=$ $\frac{c^{2}}{4\pi\hbar}\,\frac{\widetilde{\Pi}_{\mu\nu}(\omega,\mbox{$q$})}{i\omega}$ (4)
$\equiv$ $\frac{c^{2}}{4\pi\hbar}\,\frac{\Pi_{\mu\nu}(\omega,\mbox{$q$})-\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{$q$})}{i\omega}.$
Note that Eq. (4) differs from the corresponding Eqs. (3), (93), and (95) in Ref. 1 by the constant factor which is not important in the given context.
In Ref. 1 , Eq. (4) is derived from the Kubo formula expressing the conductivity tensor via the current-current correlator. In this derivation, the nonrelativistic realization of causality has been used reflected in the one-sided Fourier transformation of all relevant quantities. In the framework of the Dirac model, however, graphene is described by the relativistically covariant Dirac equation. The quantum field theoretical formalism appropriate in this case employs the Feynman Green functions, which originate from a time-ordered product of field operators and describe the scattering of particles in both directions in time. This corresponds to the relativistic causality reflected in using the two-sided Fourier transformation. In the framework of relativistic formalism appropriate for graphene described by the Dirac model, one arrives to Eq. (3) with no any “regularization” 19 ; 19a ; 19b ; 19c ; 19d .
From Eq. (2) one obtains
$\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{$q$})=\left(\begin{array}[]{ccc}\lim\limits_{\omega\to 0}\Pi_{00}&0&0\\ 0&-\frac{q_{2}^{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi&\frac{q_{1}q_{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi\\ 0&\frac{q_{1}q_{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi&-\frac{q_{1}^{2}}{q^{4}}\lim\limits_{\omega\to 0}\Pi\end{array}\right).$ (5)
Due to the gauge invariance, the polarization tensor $\Pi_{\mu\nu}$ satisfies the transversality condition $q_{\mu}\Pi^{\mu\nu}(\omega,\mbox{$q$})=0$. Due to Eq. (3), the same condition $q_{\mu}\sigma^{\mu\nu}(\omega,\mbox{$q$})=0$ is also valid for the components of the tensor of electric conductivity.
It is easily seen, however, that the expressions $\widetilde{\Pi}_{\mu\nu}$ and $\sigma_{\mu\nu}^{\rm K}$ defined in Eq. (4) do not satisfy the transversality condition. Really, from Eq. (5) one finds that
$q_{\mu}\widetilde{\Pi}^{\mu 0}(\omega,\mbox{$q$})=-q_{\mu}\lim\limits_{\omega\to 0}\Pi^{\mu 0}(\omega,\mbox{$q$})=-q_{0}\lim\limits_{\omega\to 0}\Pi^{00}(\omega,\mbox{$q$})\neq 0$ (6)
because even for the sheet of pristine graphene ($\Delta=\mu=0$) at fixed $q$ it holds 21
$\lim\limits_{\omega\to 0}\Pi^{00}(\omega,\mbox{$q$})=\frac{\alpha\pi\hbar cq}{v_{F}}+\frac{16\alpha c}{v_{F}^{2}}k_{B}T\ln 2,$ (7)
where $\alpha=e^{2}/(\hbar c)$ is the fine structure constant and $v_{F}\approx c/300$ is the Fermi velocity of graphene.
Using Eqs. (4), (6) and (7), we find
$q_{\mu}\sigma^{{\rm K},\mu 0}(\omega,\mbox{$q$})=-i\frac{\alpha c^{2}}{4\pi\hbar v_{F}}\left(\pi\hbar q+\frac{16k_{B}T}{v_{F}}\ln 2\right)\neq 0,$ (8)
i.e., for the modified expressions for the polarization and conductivity used in Ref. 1 the transversality condition for $\nu=0$ is violated. In fact the quantities $\widetilde{\Pi}_{\mu\nu}$ and $\sigma_{\mu\nu}^{\rm K}$ are not tensors because the subtraction made in Eq. (4) violated a tensorial structure.
It has been known that the i,j=1, 2i,\,j=1,\,2 components of the polarization and conductivity tensors can be presented in terms of the longitudinal and transverse quantities 22
$\Pi_{ij}(\omega,\mbox{$q$})=\frac{q_{i}q_{j}}{q^{2}}\Pi_{L}(\omega,\mbox{$q$})+\left(\delta_{ij}-\frac{q_{i}q_{j}}{q^{2}}\right)\Pi_{T}(\omega,\mbox{$q$}),$
$\sigma_{ij}(\omega,\mbox{$q$})=\frac{q_{i}q_{j}}{q^{2}}\sigma_{L}(\omega,\mbox{$q$})+\left(\delta_{ij}-\frac{q_{i}q_{j}}{q^{2}}\right)\sigma_{T}(\omega,\mbox{$q$}).$ (9)
We set the components of the first equality in Eq. (9) equal to the corresponding components in Eq. (2) and obtain
$\Pi_{L}(\omega,\mbox{$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\Pi_{00}(\omega,\mbox{$q$}),\quad\Pi_{T}(\omega,\mbox{$q$})=-\frac{1}{q^{2}}\Pi(\omega,\mbox{$q$}).$ (10)
In a similar way, using the conductivity tensor (3) and the second equality in Eq. (9), one finds
$\sigma_{L}(\omega,\mbox{$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\sigma_{00}(\omega,\mbox{$q$}),\quad\sigma_{T}(\omega,\mbox{$q$})=-\frac{1}{q^{2}}\sigma(\omega,\mbox{$q$}),$ (11)
where
$\sigma(\omega,\mbox{$q$})=q^{2}\sigma_{\mu}^{\,\mu}(\omega,\mbox{$q$})+\left(\frac{\omega^{2}}{c^{2}}-q^{2}\right)\sigma_{00}(\omega,\mbox{$q$})=\frac{c^{2}}{4\pi\hbar i\omega}\Pi(\omega,\mbox{$q$}).$ (12)
By representing the modified expressions for the polarization and conductivity (4) used in Ref. 1 in the form of Eq. (9) and repeating the same calculations with the help of Eqs. (2) and (5), we arrive at
$\widetilde{\Pi}_{L}(\omega,\mbox{$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\Pi_{00}(\omega,\mbox{$q$})=\Pi_{L}(\omega,\mbox{$q$}),$
$\widetilde{\Pi}_{T}(\omega,\mbox{$q$})=-\frac{1}{q^{2}}\left[\Pi(\omega,\mbox{$q$})-\lim\limits_{\omega\to 0}\Pi(\omega,\mbox{$q$})\right]$ (13)
and
${\sigma}_{L}^{\rm K}(\omega,\mbox{$q$})=\frac{\omega^{2}}{c^{2}q^{2}}\sigma_{00}(\omega,\mbox{$q$})=\sigma_{L}(\omega,\mbox{$q$}),$
$\sigma_{T}^{\rm K}(\omega,\mbox{$q$})=-\frac{1}{q^{2}}\left[\sigma(\omega,\mbox{$q$})-\lim\limits_{\omega\to 0}\sigma(\omega,\mbox{$q$})\right]$
[…]
As correctly concluded in Ref. 1 , the made modification (4) does not change the longitudinal conductivity of graphene and modifies only the transverse one. Below we focus on the transverse part and demonstrate that the results obtained using this modification are physically unacceptable.
Thus, for a pristine graphene at $T=0$ the quantity $\Pi$ is given by 2 ; 3 ; 4 ; 5 ; 13 ; 14
$\Pi(\omega,\mbox{$q$})=\frac{\pi e^{2}q^{2}}{c^{2}}\left\{\begin{array}[]{ll}\phantom{\mp i}\sqrt{v_{F}^{2}q^{2}-\omega^{2}},&|\omega|<v_{F}q,\\[8.53581pt] \mp i\sqrt{\omega^{2}-v_{F}^{2}q^{2}},&|\omega|>v_{F}q,\end{array}\right$. (15)
where here and below the upper and lower signs are for $\omega>v_{F}q$ and $\omega<-v_{F}q$, respectively.
From Eq. (15) at fixed $q$ one finds
$\lim\limits_{\omega\to 0}\Pi(\omega,\mbox{$q$})=\frac{\pi e^{2}v_{F}q^{3}}{c^{2}}.$ (16)
This result is valid not only at $T=0$ but at any temperature 21 . Using Eqs. (14)-(16), for the modified transverse conductivity we obtain
$\sigma_{T}^{\rm K}(\omega,\mbox{$q$})=\frac{\sigma_{0}}{\omega}\left\{\begin{array}[]{ll}i\left(\!\sqrt{v_{F}^{2}q^{2}-\omega^{2}}-v_{F}q\right),&|\omega|<v_{F}q,\\[8.53581pt] \pm\sqrt{\omega^{2}-v_{F}^{2}q^{2}}-iv_{F}q,&|\omega|>v_{F}q,\end{array}\right$. (17)
where $\sigma_{0}=e^{2}/(4\hbar)$ is the universal conductivity of graphene. Note that the regions $0<\omega<v_{F}q$ and $v_{F}q<\omega<cq$ correspond to the evanescent waves whereas for the propagating waves $0\leqslant cq\leqslant\omega$ holds.
As is seen in Eq. (17), for $\omega\geqslant cq$ Ref. 1 arrives at
${\rm Im}\,\sigma_{T}^{\rm K}(\omega,\mbox{$q$})=-\frac{\sigma_{0}v_{F}q}{\omega}$ (18)
in contradiction with the previously obtained conclusion that at zero temperature the conductivity of pure graphene in the region of propagating waves is real at all frequencies 23 ; 24 . The result following from the quantum field theoretical definition (3) of the conductivity tensor
$\sigma_{T}(\omega,\mbox{$q$})=\frac{\sigma_{0}}{\omega}\left\{\begin{array}[]{ll}i\sqrt{v_{F}^{2}q^{2}-\omega^{2}},&|\omega|<v_{F}q,\\[8.53581pt] \pm\sqrt{\omega^{2}-v_{F}^{2}q^{2}},&|\omega|>v_{F}q\end{array}\right$. (19)
is in agreement with this conclusion.
According to Ref. 1 , the relation between the polarization tensor and the electric current
$J_{\mu}(\omega,\mbox{$q$})=\frac{c}{4\pi\hbar}\Pi_{\mu\nu}(\omega,\mbox{$q$})A^{\nu}(\omega,\mbox{$q$})$ (20)
used in Refs. 10 ; 12 ; 15 ; 16 ; 19 for a derivation of Eq. (3) is unsatisfactory because it leads to a nonzero electric current expressed via $\lim\limits_{\omega\to 0}\Pi_{\mu\nu}(\omega,\mbox{$q$})$ for the equal to zero electric field. Based on this, Ref. 1 replaces $\Pi_{\mu\nu}$ in Eq. (20) with the modified expression $\widetilde{\Pi}_{\mu\nu}$ defined in Eq. (4). Below we show that this replacement has no justification.
[…]
$\lim\limits_{\omega\to 0}\Pi_{ii}(\omega,\mbox{$q$})=-\left(1-\frac{q_{i}^{2}}{q^{2}}\right)\frac{\pi e^{2}v_{F}q}{c^{2}}\neq 0.$ (21)
Thus, in the region of evanescent waves, at any temperature, including $T=0$, the transverse conductivity $\sigma_{T}$ found in the framework of quantum field theory has an imaginary part. For instance, at $T=0$, where it is pure imaginary, using the first line of Eq. (15), we obtain
$\sigma_{T}(\omega,\mbox{$q$})=i\sigma_{0}\frac{\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}{\omega}.$ (22)
According to Ref. 1 , Eq. (21) results in the appearance of an electric current in graphene in the absence of electric field, whereas Eq. (22) leads to an unacceptable double pole at zero frequency in the graphene dielectric permittivity. If the expression (4) is used 1 , at zero temperature the modified polarization, $\widetilde{\Pi}_{\mu\nu}(\omega,\mbox{$q$})$, and conductivity, $\sigma_{T}^{\rm K}(\omega,\mbox{$q$})$, expressions vanish in the limit of zero frequency.
Note, however, that at nonzero temperature in the region of propagating waves the pure imaginary current in the absence of electric field arises both in the quantum field theoretical formalism using Eq. (20) and in the approach of Ref. 1 which replaces the polarization tensor $\Pi_{\mu\nu}$ in Eq. (20) with the modified expression, $\widetilde{\Pi}_{\mu\nu}$, defined in Eq. (4).
Really, for the propagating waves the condition $0\leqslant cq\leqslant\omega$ holds. Then, if $\omega\to 0$, the wave vector $q$ must go to zero as well. For the real parts of $\Pi_{00}$ and $\Pi$, under the condition $\hbar\omega\ll k_{B}T$, one has 2
${\rm Re}\,\Pi_{00}(\omega,\mbox{$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar}\,\frac{q^{2}}{\omega^{2}}\left[1+O\left(\frac{v_{F}^{2}q^{2}}{\omega^{2}}\right)\right],$
${\rm Re}\,\Pi(\omega,\mbox{$q$})=8\ln 2\frac{e^{2}k_{B}T}{\hbar}\,\frac{q^{2}}{c^{2}}\left[1+O\left(\frac{v_{F}^{2}q^{2}}{\omega^{2}}\right)\right].$ (23)
With the help of these results, using Eq. (2), we find
${\rm Re}\,\Pi_{11}(\omega,\mbox{$q$})={\rm Re}\,\Pi_{22}(\omega,\mbox{$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}},$ (24)
i.e., these components of the polarization tensor in the lowest order with respect to the small parameter $v_{F}^{2}q^{2}/\omega^{2}$ are equal to the constant independent on $\omega$ and $q$. Thus, they preserve their value (24) in the limit $\omega,\,q\to 0$.
Using Eqs. (10) and (23), we also obtain
${\rm Re}\,\Pi_{L}(\omega,\mbox{$q$})={\rm Re}\,\Pi_{T}(\omega,\mbox{$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}.$ (25)
The corresponding values of the longitudinal and transverse conductivities are obtained from Eqs. (3), (11), (12), and (23)
${\rm Im}\,\sigma_{L}(\omega,\mbox{$q$})={\rm Im}\,\sigma_{T}(\omega,\mbox{$q$})=\sigma_{0}\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega}.$ (26)
Note that in the region of propagating waves the result (26) leads to a simple pole in the corresponding dielectric permittivities of graphene
${\rm Re}\,\varepsilon_{{\rm L,T}}(\omega,\mbox{$q$})=1-\frac{2\pi q}{\omega}{\rm Im}\,\sigma_{{\rm L,T}}(\omega,\mbox{$q$})$
$~~=1-16\ln 2\frac{\sigma_{0}}{c}\,\sin\theta\,\frac{k_{B}T}{\hbar\omega},$ (27)
where $\theta$ is the angle of incidence of the wave on a graphene sheet, $\sin\theta=cq/\omega=\mbox{const}$.
For the modified expressions used in Ref. 1 from Eqs. (13), (16), and (23) one finds
${\rm Re}\,\widetilde{\Pi}_{L}(\omega,\mbox{$q$})={\rm Re}\,\Pi_{L}(\omega,\mbox{$q$}),$
${\rm Re}\,\widetilde{\Pi}_{T}(\omega,\mbox{$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}\,+\frac{\pi e^{2}v_{F}q}{c^{2}}.$ (28)
The corresponding conductivities found from Eq. (14) have the imaginary parts
${\rm Im}\,\sigma_{L}^{\rm K}(\omega,\mbox{$q$})=\sigma_{0}\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega},$
${\rm Im}\,\sigma_{T}^{\rm K}(\omega,\mbox{$q$})=\sigma_{0}\left(\frac{8\ln 2}{\pi}\,\frac{k_{B}T}{\hbar\omega}-\frac{v_{F}q}{\omega}\right).$ (29)
What is important, the diagonal components of the modified polarization expression found from Eq. (9) with added tildes using Eq. (28) do no vanish in the limit $\omega,\,q\to 0$
$\lim\limits_{\omega,\,q\to 0}{\rm Re}\,\widetilde{\Pi}_{11}(\omega,\mbox{$q$})=\lim\limits_{\omega,\,q\to 0}{\rm Re}\,\widetilde{\Pi}_{22}(\omega,\mbox{$q$})=-8\ln 2\frac{e^{2}k_{B}T}{\hbar c^{2}}.$ (30)
Thus, at nonzero temperature in both the quantum field theory and in the formalism of Ref. 1 a nonzero current in graphene arises even for a zero electric field. Keeping in mind, however, that the conductivity in both cases is pure imaginary because its real part vanishes when $\omega,\,q\to 0$ 2 , this creates no problem. Such a pure imaginary current satisfies the relativistically covariant Ohm’s law 24a
$J_{\mu}(\omega,\mbox{$q$})=\sigma_{\mu\nu}(\omega,\mbox{$q$})\,E^{\nu},$ (31)
where $\sigma_{\mu\nu}$ goes to $i\infty$ with vanishing $E^{\nu}$. Note that according to Ref. 1 the current in Eq. (31) is different from that in Eq. (20) because the latter contains the magnetically induced part. This statement is, however, incorrect. In the absence of the constant in time, external magnetic field, both Eqs. (20) and (31) describe the total induced current 19a .
It seems illogical that the same behavior of $\sigma_{L}^{\rm K}$ and $\sigma_{T}^{\rm K}$ in Eq. (29) following from the modified formalism of Ref. 1 in the region of propagating waves as of $\sigma_{L}$ and $\sigma_{T}$ in Eqs. (22) and (26) using quantum field theory is not considered in Ref. 1 as leading to a permanent electric current in the absence of electric field. It should be noted that Ref. 1 introduces the phenomenological parameter $\Gamma$ describing the dissipation of electronic quasiparticles by replacing the frequency $\omega$ with $\omega+i\Gamma$ (see below). At the cost of this replacement, the pure imaginary currents arising in the absence of nonzero electric field can be eliminated. However, an incorporation of the dissipation of electronic quasiparticles brings the approach of Ref. 1 outside the application region of the Dirac model, where the quasiparticles are considered as noninteracting. Note also that just the behavior of the conductivities of graphene according to Eq. (26) leads to the big thermal effect in the Casimir force between two graphene sheets at short separations predicted in Ref. 25 for the case of two pristine graphene sheets and confirmed experimentally in Refs. 26 ; 27 .
Hence, the modified expression (4) was obtained in Ref. 1 by inappropriately using the nonrelativistic realization of the causality condition in the relativistic Dirac model. This modification caused a violation of the condition of gauge invariance and the physical inconsistencies described above. As to a prediction of the double pole at zero frequency by quantum field theory, it does not contradict to any physical results, including the Kubo approach formulated for the propagating fields, and suggests a path to the resolution of long-term problems in the Casimir effect 13 .
[…]
What is more, the expression connecting the polarization and conductivity tensors is written in the form used in Refs. 10 ; 12 where the system of units with $\hbar=c=k_{B}=\varepsilon_{0}=1$ has been used ($\varepsilon_{0}$ is the dielectric permittivity of vacuum). In contradiction with this, Ref. 1 uses the expression for the fine structure constant $\alpha=e^{2}/(\hbar c)$, whereas in Refs. 10 ; 12 the expression $\alpha=e^{2}/(4\pi)$ has been used. The definition of $q_{z}$ below Eq. (120) is in contradiction with the definition of the same quantity given in TABLE I, right column, line 9. Above Eq. (C34) Ref. 1 considers “the case $\Pi_{\mu\nu}(q)=\Pi_{L}(q)$” which is meaningless if to take into account the expression for $\Pi_{L}(q)$ in Eq. (121).
To conclude, in the foregoing it was shown that the modification of the polarization and conductivity tensors made in Ref. 1 violates their transversality and leads to other physically unacceptable consequences. This modification is superfluous because the claims of Ref. 1 about the problems arising when using the standard formalism of quantum field theory are invalid. Equation (4) was derived from the Kubo formula by using the nonrelativistic realization of causality in application to the relativistic Dirac model where it is inapplicable.
It should be added also that, according to Ref. 1 , the quantum field theoretical description of the conductivity of graphene does not describe unavoidable losses. This is, however, not the case because the conductivity of graphene obtained using this description possesses both the real and imaginary parts. In doing so, the real part of conductivity results in the positive imaginary part of the dielectric permittivity of graphene which describes the losses of conduction electrons in graphene. As to the massless or very light electronic quasiparticles, in the application region of the Dirac model they are considered as noninteracting. By introducing the phenomenological parameter $\Gamma$ for a description the relaxation of quasiparticles in graphene, Ref. 1 again violated the relativistic description of graphene inherent to the Dirac model.
All the above examples were given for the case of a pristine graphene in order to avoid unnecessary technical difficulties and to make the presentation maximally transparent. However, in the application region of Dirac model, all the statements made remain valid for a general case of graphene possessing some nonzero energy gap and chemical potential.
The authors are grateful to D. V. Vassilevich for useful comments. The work of G.L.K. and V.M.M. was supported by the State Assignment for Basic Research (project FSEG-2023-0016). N.K. was supported in part by the grants 2025/13673-0, 2021/10128-0 of São Paulo Research Foundation (FAPESP).